# Interpretable Scaling Behavior in Sparse Subnetwork Representations of Quantum States

### v1

@@ 4.2 Sparse scaling behavior
In accordance with pmlr-v139-rosenfeld21a , the scaling behavior of the pruned FFNN error shown in Fig. 1 suggests three scaling regimes: (i) a low-error plateau (LEP) at high densities $n$ (ii) a high-error plateau (HEP) at low densities and (iii) an approximate power-law region (APL) connecting the two plateaus. We find the error scaling in the (HEP) and the (APL) to be universal with respect to the width of the initial network; Fig. 2 demonstrates evidence that the error of all shallow FFNNs of different width coalesce onto one another. However, the LEP at large number of remaining parameters shows a slight dependence on the initial number of network parameters: These denser networks are qualitatively different than the very sparse networks in the APL and HEP, involving a more complex structure that is broken within the APL. We will further elaborate on this question in Section 4.3.
@@ 4.3 Sparsity-induced phase transition
What do we know about the quantum phase of $\Psi^{NN}(\rho)$ a priori? When $\rho$ is very large and the neural-network approximation is very good the quantum phase of $\Psi^{NN}(\rho)$ will correspond to the same phase as the targeted state $\Psi^{exact}_{\kappa=\kappa_{c}}$, i.e. the exact ground state of the TFIM at $\kappa=\kappa_{c}$. We expect this to be true in the LEP. However, $\Psi^{NN}(\rho)$ may undergo a phase transition to a phase that is easier to represent as $\rho$ is decreased. We first found evidence of this behavior in a kink of the network error at the transition between the APL and the HEP. This signifies a discontinuity in $\frac{dE}{d\rho}$, which is a typical signature of a (first-order) phase transition.
@@ 4.4 Interpretable sparse neural network solution to the toric code
This asymptotically exact solution consists of plaquette-shaped filters with three positive weights and one negative weight (each of magnitude $W$) or vice-versa as shown in Fig. 4, which we call "odd parity filters". We can interpret this structure by considering an exact toric code ground state solution, which consists of an equal superposition of all spin-configurations $|\bm{\sigma}\rangle$ that fulfill the “plaquette conditions” $\lambda_{B_{p}}(\bm{\sigma})=\prod_{i\in p}\sigma^{z}_{i}=1$ $\forall p$ kitaev2003fault . In other words, to represent this ground state a neural-network solution has to fulfill $\Psi^{NN}(\bm{\sigma})=1$ if $\lambda_{B_{p}}(\bm{\sigma})=1$ $\forall p$, and $\Psi^{NN}(\bm{\sigma})=0$ else. Then, together with its activation function, an odd parity filter connected to plaquette $p$ has the effect of exponentially amplifying the amplitude $\Psi^{NN}(\bm{\sigma})$ for spin-configurations $\bm{\sigma}$ that fulfill $\lambda_{B_{p}}(\bm{\sigma})=1$ relative to configurations with $\lambda_{B_{p}}(\bm{\sigma})=-1$ (see Table 1). In total, the excited state contributions are suppressed relative to the ground state by a factor of $\exp(4W)$ for every broken plaquette condition. In the limit of $W\to\infty$, we thus arrive at the exact solution.Further details on this solution can be found in the Appendix.
@@
(4)(4$\uparrow$)( 4 ↑ ), (2,2)(2$\uparrow$,2$\downarrow$)( 2 ↑ , 2 ↓ ), (4)(4$\downarrow$)( 4 ↓ ) yes $W$ $\exp(8W)$
[…]
@@ 5 Conclusion
Finally, we show that through pruning we are able to discover a new (asymptotically exact) solution to the toric code. This demonstrates that pruning neural networks can elucidate how they are able to solve problems.
Understanding how to obtain a neural-network quantum state representation with minimal number of parameters may in addition be relevant for applications such as the time evolution of many-body quantum states: A main computational cost here typically lies in inverting an $m\times m$ matrix where $m$ is the number of network parameters. We leave the exploration of this idea to future studies.
@@
The code, all pretrained models, and data are available on Zenodo anonymous_2025_15492729 .
@@ Appendix A Pruning procedures
  for each pruning iteration in $I$ do
     Select $p_{\mathrm{r}}$of weights (among unmasked) with lowest magnitude.
@@ Appendix B Scaling behavior in the variance of the energy
Appendix C Sparse neural network solution to the toric code
@@
Appendix D Neural network architectures
@@
Appendix E Sampling and optimization procedures
@@
where the local energy is defined by Eq. 18. To ensure that the inverse of the S-matrix in Eq. 20 is well defined, we use an explicit regularization $S=S+\lambda I$, where $\lambda$ is called the diagonal shift (DS). We keep the learning rate and diagonal shift constant throughout this work, and report the values we use in Section F.
@@
Appendix F Resources for reproducability
@@
$I$ Total number of pruning iterations All
@@
All symbols, acronyms, and hyperparameters used in this work are reported in Table 2. We report the all hyperparameters and training settings used across all experiments in Tables 3, 4, 5, and 6 to ensure reproducibility and facilitate comparison in future work.
@@
Network $w/n_{f}$ $d/n_{b}$ $k_{d}$ $j$ $k$ $p_{r}$ $I$ $n_{\mathrm{init}}$ $N_{\mathrm{s}}$ $\eta$ $\lambda$
@@
Panel $\kappa$ $w$ $d$ $j$ $k$ $p_{r}$ $I$ $n_{\mathrm{init}}$ $N_{\mathrm{s}}$ $\eta$ $\lambda$
(a) 3.04438 $N$ 1 1e4 1e3 0.12 54 10000 1024 8e-3 1e-4
@@
System size $N$ $w$ $d$ $j$ $k$ $p_{r}$ $I$ $n_{\mathrm{init}}$ $N_{\mathrm{s}}$ $\eta$ $\lambda$
@@
Panel(s) $w$ $d$ $j$ $k$ $p_{r}$ $I$ $n_{\mathrm{init}}$ $N_{\mathrm{s}}$ $\eta$ $\lambda$

### v2

@@ 4.2 Sparse scaling behavior
In accordance with Rosenfeld et al. (2021), the scaling behavior of the pruned FFNN error shown in Fig. 1 suggests three scaling regimes: (i) a low-error plateau (LEP) at high densities $n$ (ii) a high-error plateau (HEP) at low densities and (iii) an approximate power-law region (APL) connecting the two plateaus. We find the error scaling in the (HEP) and the (APL) to be universal with respect to the width of the initial network; Fig. 2 demonstrates evidence that the error of all shallow FFNNs of different width coalesce onto one another. However, the LEP at large number of remaining parameters shows a slight dependence on the initial number of network parameters: These denser networks are qualitatively different than the very sparse networks in the APL and HEP, involving a more complex structure that is broken within the APL. We will further elaborate on this question in Section 4.3.
@@ 4.3 Sparsity-induced phase transition
What do we know about the quantum phase of $\Psi^{NN}(\rho)$ a priori? When $\rho$ is very large and the neural-network approximation is very good the quantum phase of $\Psi^{NN}(\rho)$ will correspond to the same phase as the targeted state $\Psi^{exact}_{\kappa=\kappa_{c}}$, i.e. the exact ground state of the TFIM at $\kappa=\kappa_{c}$. We expect this to be true in the LEP. However, $\Psi^{NN}(\rho)$ may undergo a phase transition to a phase that is easier to represent as $\rho$ is decreased. We find evidence of this behavior in a kink of the network error at the transition between the APL and the HEP. This signifies a discontinuity in $\frac{dE}{d\rho}$, which is a typical signature of a (first-order) phase transition.
@@ 4.4 Interpretable sparse neural network solution to the toric code
This asymptotically exact solution consists of plaquette-shaped filters with three positive weights and one negative weight (each of magnitude $W$) or vice-versa as shown in Fig. 4, which we call "odd parity filters". We can interpret this structure by considering an exact toric code ground state solution, which consists of an equal superposition of all spin-configurations $|\bm{\sigma}\rangle$ that fulfill the “plaquette conditions” $\lambda_{B_{p}}(\bm{\sigma})=\prod_{i\in p}\sigma^{z}_{i}=1$ $\forall p$ Kitaev (2003). In other words, to represent this ground state a neural-network solution has to fulfill $\Psi^{NN}(\bm{\sigma})=1$ if $\lambda_{B_{p}}(\bm{\sigma})=1$ $\forall p$, and $\Psi^{NN}(\bm{\sigma})=0$ else. Then, together with its activation function, an odd parity filter connected to plaquette $p$ has the effect of exponentially amplifying the amplitude $\Psi^{NN}(\bm{\sigma})$ for spin-configurations $\bm{\sigma}$ that fulfill $\lambda_{B_{p}}(\bm{\sigma})=1$ relative to configurations with $\lambda_{B_{p}}(\bm{\sigma})=-1$ (see Table 1). In total, the excited state contributions are suppressed relative to the ground state by a factor of $\exp(4W)$ for every broken plaquette condition. In the limit of $W\to\infty$, we thus arrive at the exact solution. Further details on the convergence of this solution over increasing magnitude weights can be found in the Appendix.
@@
$(4\uparrow)$, $(2\uparrow,2\downarrow)$, $(4\downarrow)$ $W$ $\exp(8W)$
$(3\uparrow,1\downarrow)$, $(1\uparrow,3\downarrow)$ 0.5WW $\exp(4W)$
@@ 5 Conclusion
Finally, we show that through pruning we are able to discover a new (asymptotically exact) solution to the toric code which is simpler than human-conceptualized solutions Carrasquilla and Melko (2017); Chen et al. (2025b); Deng et al. (2017); Valenti et al. (2022), and does not require a priori architecture design principles. This demonstrates that pruning neural networks can elucidate how they are able to solve problems. We expect that pruning may uncover further exact solutions to other Hamiltonians, including those with topologically ordered ground states, those representing quantum error correction codes.
Understanding how to obtain a neural-network quantum state representation with minimal number of parameters may be relevant when it is beneficial to have a compressed model for further processing. For example, time evolution Sinibaldi et al. (2023), quantum circuit simulation Medvidović and Carleo (2021), and fine-tuning neural network quantum states Rende et al. (2024) are all applications where an accurate compressed model could be advantageous. Furthermore, sparse network representations of quantum states may unlock the capability to use desired optimization strategies. Equipped with a pruned neural network that represents the initial state accurately, the fact that the architecture utilizes a small amount of parameters may allow for the use of SR in the $n<s$ regime where $n$ the number of parameters and $s$ is the number of samples. The ability to use SR may provide benefits over minSR Chen and Heyl (2024), a variant of SR which is more efficient in the $n>s$ regime, but limited in terms of the number of samples one can use in practice. We leave the exploration of this idea to future studies.
@@
The code, all pretrained models, and data are available on Zenodo Anonymous (2025).
@@ Appendix A Pruning procedures
for each pruning iteration in II do
  Select $p_{\mathrm{r}}$% of weights (among unmasked) with lowest magnitude.
@@ Appendix B Scaling behavior in the variance of the energy
Appendix C Robustness over many random initializations
Table 2: Robustness of $m_{\mathrm{imp}}$ mask structure at different iterations of pruning for the $N=10\times 10$ TFIM (corresponding to panel (b) of Fig. 1), and $N=18$ toric code (corresponding to panel (a) of Fig. 4) Hamiltonians. We report the original relative energy error for $\theta_{\mathrm{init}}$ and the mean relative energy error over 5 different initializations for $\theta_{\mathrm{rand}}$, in addition to their 99.9% confidence interval.
Hamiltonian Pruning iteration Initialization type
$\theta_{\mathrm{init}}$ $\theta_{\mathrm{rand}}$ 99.9% CI
TFIM $i=12$ 1.54e-4 1.50e-4 [1.44e-4, 1.56e-4]
$i=15$ 1.65e-4 1.67e-4 [1.58e-4, 1.77e-4]
$i=18$ 1.96e-4 1.99e-4 [1.50e-4, 2.19e-4]
TC $i=9$ 2.94e-5 1.83e-5 [0, 4.55e-5]
$i=12$ 1.64e-5 1.52e-5 [0, 3.45e-5]
$i=15$ 9.71e-6 1.60e-5 [0, 4.30e-5]
To substantiate the claim of the importance of the structure in the masks, and the relative unimportance of parameter initialization, we test the parameter sensitivity in the reduced model dimension for the $T(\theta_{\mathrm{rand}},m_{\mathrm{imp}})$ isolated training tests, starting from many random initializations. We confirm the robustness of the subnetwork structure by reporting the mean of five additional random instances of the isolated training variant at several pruning iterations and their 99.9% confidence interval shown in Table 2. All random initializations remain close to the $\theta_{\mathrm{init}}$ initialization and are in agreement at the 99.9% level. This confirms the robustness of the mask structure, and demonstrates that under many random initializations, a good mask is sufficient to obtain a matching error to the lottery ticket initialization.
Appendix D Sparse neural network solution to the toric code
@@
Appendix E Neural network architectures
@@
Appendix F Sampling and optimization procedures
@@
where the local energy is defined by Eq. 18. To ensure that the inverse of the S-matrix in Eq. 20 is well defined, we use an explicit regularization $S=S+\lambda I$, where $\lambda$ is called the diagonal shift (DS). We keep the learning rate and diagonal shift constant throughout this work, and report the values we use in Section G.
@@
Appendix G Resources for reproducability
@@
II Total number of pruning iterations All
@@
All symbols, acronyms, and hyperparameters used in this work are reported in Table 3. We report the all hyperparameters and training settings used across all experiments in Tables 4, 5, 6, and 7 to ensure reproducibility and facilitate comparison in future work.
@@
Network $w/n_{f}$ $d/n_{b}$ $k_{d}$ $j$ $k$ $p_{r}$ II $n_{\mathrm{init}}$ $N_{\mathrm{s}}$ $\eta$ $\lambda$
@@
Panel $\kappa$ $w$ $d$ $j$ $k$ $p_{r}$ II $n_{\mathrm{init}}$ $N_{\mathrm{s}}$ $\eta$ $\lambda$
(a) 3.04438 NN 1 1e4 1e3 0.12 54 10000 1024 8e-3 1e-4
@@
System size NN $w$ $d$ $j$ $k$ $p_{r}$ II $n_{\mathrm{init}}$ $N_{\mathrm{s}}$ $\eta$ $\lambda$
@@
Panel(s) $w$ $d$ $j$ $k$ $p_{r}$ II $n_{\mathrm{init}}$ $N_{\mathrm{s}}$ $\eta$ $\lambda$